### v1

@@ Abstract
Spatial transformations that capture population-level morphological statistics are critical for medical image analysis. Commonly used smoothness regularizers for image registration fail to integrate population statistics, leading to anatomically inconsistent transformations. Inverse consistency regularizers promote geometric consistency but lack population morphometrics integration. Regularizers that constrain deformation to low-dimensional manifold methods address this. However, they prioritize reconstruction over interpretability and neglect diffeomorphic properties, such as group composition and inverse consistency. We introduce MORPH-LER, a Log-Euclidean regularization framework for population-aware unsupervised image registration. MORPH-LER learns population morphometrics from spatial transformations to guide and regularize registration networks, ensuring anatomically plausible deformations. It features a bottleneck autoencoder that computes the principal logarithm of deformation fields via iterative square-root predictions. It creates a linearized latent space that respects diffeomorphic properties and enforces inverse consistency. By integrating a registration network with a diffeomorphic autoencoder, MORPH-LER produces smooth, meaningful deformation fields. The framework offers two main contributions: (1) a data-driven regularization strategy that incorporates population-level anatomical statistics to enhance transformation validity and (2) a linearized latent space that enables compact and interpretable deformation fields for efficient population morphometrics analysis. We validate MORPH-LER across two families of deep learning-based registration networks, demonstrating its ability to produce anatomically accurate, computationally efficient, and statistically meaningful transformations on the OASIS-1 brain imaging dataset. Github: https://github.com/iyerkrithika21/MORPH_LER
@@ 1 Introduction
Deep learning methods address these challenges by learning complex transformations directly from images, offering faster inference and ensuring smooth deformation fields through regularizers like bending energy or $L_{2}-$gradient penalties [Balakrishnan et al.(2019)Balakrishnan, Zhao, Sabuncu, Guttag, and Dalca, Chen et al.(2022)Chen, Frey, He, Segars, Li, and Du]. However, they treat registration as a computational problem, neglecting diffeomorphic properties like inversion and composition, resulting in biologically implausible solutions. Improvements like gradient and transformation inverse consistency enhance diffeomorphic transformations [Tian et al.(2023)Tian, Greer, Vialard, Kwitt, Estépar, Rushmore, Makris, Bouix, and Niethammer]. Yet, existing approaches often predict transformations without deeper insights into underlying anatomical relationships, oversimplifying interpretations. Nevertheless, the mathematical framework of diffeomorphic transformations, defined by the diffeomorphism group $Diff(\mathcal{M})$ ($dim(\mathcal{M})>0$) which is a Fréchet Lie group and manifold, provides a rigorous foundation for capturing and analyzing complex anatomical variability. Unlike Euclidean spaces, operations like addition or averaging cannot be defined in $Diff(\mathcal{M})$. Techniques such as Principal Geodesic Analysis (PGA) [Fletcher et al.(2004)Fletcher, Lu, Pizer, and Joshi], Fréchet means [Le and Kume(2000)], and geodesic regression [Fletcher(2011)] attempt to address statistical estimation for Riemannian manifolds but require significant adaptation for Frechet Lie groups. Therefore, alternative metrics are necessary to model relationships between deformation fields and ensure anatomical relevance. The Log-Euclidean framework [Arsigny et al.(2006)Arsigny, Commowick, Pennec, and Ayache] transforms non-linear diffeomorphism manifolds into linear Lie algebra spaces, enabling efficient computation of transformation distances and statistical analysis. However, iterative processes used in logarithm estimation can be susceptible to initializations, potentially leading to suboptimal results. Standalone deformation analysis models address the limitations of the iterative methods and provide a computationally efficient transformation of diffeomorphic groups into linear Lie algebra. These mathematical foundations provide a rigorous framework for analyzing the complex structure of diffeomorphisms, leading to more reliable and interpretable results.
@@ 2 Related Work
Population-aware registration networks, such as CAE [Bhalodia et al.(2019)Bhalodia, Elhabian, Kavan, and Whitaker], add regularization by limiting deformation fields to low-dimensional manifolds, effectively capturing anatomical variability. However, they often fall short of explicitly enforcing diffeomorphism essential for interpretability and robustness. CAE’s regularization approach begins with an L2 regularizer, providing initial stability. However, when the autoencoder-based regularizer is introduced, it struggles to preserve anatomical consistency, often resulting in non-diffeomorphic transformations.
Similarly, diffeomorphic autoencoders for LDDMM [Hinkle et al.(2018)Hinkle, Womble, and Yoon] have been proposed for atlas building by encoding velocity field-based transformations into latent spaces for statistical analysis. These methods face computational challenges with large datasets and complex anatomical variations. Additionally, the focus on deformation field reconstruction from latent space representation often sacrifices interpretability, limiting their utility for downstream tasks such as population analysis. The Log Euclidean Diffeomorphic Autoencoder (LEDA) [Iyer et al.(2024)Iyer, Elhabian, and Joshi] framework has combined the advantages of the Log Euclidean framework with a deep learning approach. LEDA estimates the principal logarithms of the deformation field while maintaining inverse consistency in both the deformation and latent spaces. This approach efficiently captures complex anatomical variations while remaining computationally tractable for large-scale analysis. It maps composition in the data space to linearlized latent space, enabling vector-based statistics and ensuring robustness while adhering to diffeomorphism group laws.
@@
Figure 1: Proposed Architecture for MORPH-LER on the left and the estimated atlas using the algorithm described in section 4 compared with pixel-wise average atlas on the right
@@ 3 Methods
Background: A diffeomorphism ${\boldsymbol{\phi}}:\mathbb{R}^{D}\to\mathbb{R}^{D}$, is a smooth, invertible mapping with a smooth inverse, defined as ${\boldsymbol{\phi}}(\mathbf{x})=\mathbf{x}+\mathbf{u}(\mathbf{x})$, where $\mathbf{u}(\mathbf{x})$ is the displacement field. Diffeomorphic transformations can be constructed by composing multiple small deformations: ${\boldsymbol{\phi}}(\mathbf{x})=(\mathbf{x}+\epsilon\mathbf{v}_{1}(\mathbf{x}))\circ(\mathbf{x}+\epsilon\mathbf{v}_{2}(\mathbf{x}))\circ\cdots\circ(\mathbf{x}+\epsilon\mathbf{v}_{n}(\mathbf{x}))$. The space of diffeomorphisms $\mathrm{Diff}(\mathcal{M})$, forms a Lie group under composition, with the identity map, serving as the group’s identity element; i.e., ${\boldsymbol{\phi}}\circ\mathbf{id}=\mathbf{id}\circ{\boldsymbol{\phi}}={\boldsymbol{\phi}},\forall{\boldsymbol{\phi}}\in\mathrm{Diff}(\mathcal{M})$. The associated Lie algebra consists of smooth vector fields $\mathbf{v}(\mathbf{x}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, which generate diffeomorphisms through their flows, described by one-parameter subgroups $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$. The exponential map $\exp(\mathbf{v})$ connects the Lie algebra to the Lie group, while the logarithm map $\log({\boldsymbol{\phi}})$ serves as its local inverse. To compute the logarithm map, a non-linear inverse scaling and square rooting algorithm [Arsigny et al.(2006)Arsigny, Commowick, Pennec, and Ayache] is used:
@@
Table 1: Registration Performance Metrics: Top-performing model and next-best model
@@
Figure 2: (a) MORPH-LER Square Root Estimation (b) Validation of Small Deformation Field Assumption (c) Validation of Latent Inverse Consistency
@@ 4 Results
Figure 3: (A) MORPH-LER PCA Modes of Logarithm Maps and Latent Representations. (B) Qualitative Results for Registration. Top-performing model (bold) and next-best model (underlined) for the illustrated sample are highlighted.
@@
We use 2D coronal slices from the OASIS-1 dataset [Marcus et al.(2007)Marcus, Wang, Parker, Csernansky, Morris, and Buckner], which includes brain MRIs from 100 subjects (60 with Alzheimer’s). To evaluate our proposed regularizer, we compare it against baselines: CAE [Bhalodia et al.(2019)Bhalodia, Elhabian, Kavan, and Whitaker], an autoencoder-based regularization method, and the original GradICON [Tian et al.(2023)Tian, Greer, Vialard, Kwitt, Estépar, Rushmore, Makris, Bouix, and Niethammer] (convolutional) and TransMorph [Chen et al.(2022)Chen, Frey, He, Segars, Li, and Du] (transformer-based) without LEDA. We utilize GradICON and TransMorph as primary registration networks with LEDA as the secondary network, naming them M-GradICON and M-TransMorph.
@@
Table 1 compares TransMorph vs. M-TransMorph, GradICON vs. M-GradICON, and CAE as a standalone regularizer. Within each subgroup, we analyze trade-offs between registration accuracy and topological preservation. M-TransMorph improves diffeomorphic properties by reducing negative Jacobian pixels but slightly impacts segmentation performance. M-GradICON achieves higher Dice scores than GradICON while maintaining low negative Jacobian pixels, balancing accuracy and topology preservation. Compared to CAE, LEDA-based regularization maintains similar Dice scores while significantly reducing negative Jacobian pixels, demonstrating its superiority in preserving anatomical topology without compromising accuracy. The qualitative results presented in Figure 3.B shows the registration performance across different models for four distinct anatomical structures. These results corroborate the quantitative findings. Notably, our proposed M-GradICON model significantly improves over its baseline counterpart. This underscores the effectiveness of the secondary network. The performance of M-GradICON is especially noticeable in the smaller anatomical regions such as Subcortical-Gray-Matter and Cerebrospinal fluid (CSF). The quantitative and qualitative results (Figure 3 and Figure 4) show that the CAE regularization strategy with a simple autoencoder is ineffective for complex shapes, lacks diffeomorphic properties, and produces anatomically inconsistent transformations.
@@
Figure 2.A illustrates the square root estimations $\phi^{-m}$ of MORPH-LER variants and demonstrates a progressive warping of the source image to align with the target. M-GradICON exhibits superior performance over M-TransMorph, featuring smoother deformation grids, better alignment of warped images, and smoother Jacobian maps, indicating diffeomorphic superiority. To validate the small deformation field assumption [^1] Small deformation field assumption $u_{AB}(\mathbf{x})=u_{BA}(\mathbf{x})$ reflects symmetry of infinitesimal displacements, ensuring consistent and reversible transformations., we tested the logarithm map consistency by utilizing the $2^{6}$-th root estimation, we systematically negated and composed forward and inverse displacement fields. Figure 2.C demonstrates that both methods satisfy this fundamental assumption, validating the accuracy of logarithm maps. Furthermore, we extended this validation to the model’s latent space by negating the latent representations of forward and inverse displacement fields. As illustrated in Figure 2.B, the MORPH-LER accurately decodes these negated latent representations into their corresponding inverse fields, providing compelling evidence of inverse consistency in the latent space.
@@
PCA analysis of the MORPH-LER’s latent space, depicted in Figure 3.A reveals modes of variation that align closely with logarithm map PCA results. These modes capture clinically consistent changes, such as ventricular expansion and hippocampus atrophy, while maintaining smooth, structured transitions. This latent space representation not only supports accurate reconstruction of deformations but also enables intuitive exploration of clinically meaningful variations, highlighting the MORPH-LER’s potential for generating realistic deformations, interpolating between anatomical states, and providing valuable insights for understanding disease progression.
@@
We propose an efficient atlas estimation approach that leverages the trained MORPH-LER’s linearized latent space, which adheres to Lie group action laws. The algorithm begins by randomly selecting an initial image $\mathbf{A}^{(0)}$ as the starting atlas. For each image $\mathbf{I}_{i}$ in the dataset, the algorithm computes bidirectional transformations (${\boldsymbol{\phi}}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$ and ${\boldsymbol{\phi}}_{\mathbf{I}_{i}\mathbf{A}^{(k)}}$) between the current atlas $\mathbf{A}^{(k)}$ and the image $\mathbf{I}_{i}$. The latent representations of these transformations, $\mathbf{z}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$ and $\mathbf{z}_{\mathbf{I}_{i}\mathbf{A}^{(k)}}$, are extracted using the LEDA module. At each iteration, the latent representations corresponding to the transformations from the atlas to all images, $\mathbf{z}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$, are averaged across all $i$ to compute a mean representation, $\mathbf{z}^{k}$. This negated mean representation is decoded to obtain the atlas to image mean deformation field $\overline{{\boldsymbol{\phi}}}^{k}$, which is then used update the atlas $\mathbf{A}^{(k)}$ pulling it towards the true mean. The process is repeated until the atlas converges to a stable solution, as determined by minimal changes across iterations. Additional details are provided in Appendix.
To evaluate the robustness of this method, we perform multiple estimations, each initialized with a different atlas. The initial atlases are chosen from two distinct age groups: above 45 years and 45 years or below. As shown in Figure 1, the final estimated atlas remains consistent regardless of the initialization, demonstrating the stability of the approach. The atlas serves as a population-level representation that captures shared anatomical features while preserving structural details. Unlike a naive pixel-wise average, it reflects consistent geometric alignment across the dataset, making it critical for studying population variability and facilitating downstream statistical analysis. This ensures the atlas is both biologically meaningful and robust for comparative studies.

### v2

@@ Abstract
Spatial transformations that capture population-level morphological statistics are critical for medical image analysis. Commonly used smoothness regularizers for image registration fail to integrate population statistics, leading to anatomically inconsistent transformations. Inverse consistency regularizers promote geometric consistency but lack population morphometrics integration. Regularizers that constrain deformation to low-dimensional manifold methods address this. However, they prioritize reconstruction over interpretability and neglect diffeomorphic properties, such as group composition and inverse consistency. We introduce MORPH-LER, a Log-Euclidean regularization framework for population-aware unsupervised image registration. MORPH-LER learns population morphometrics from spatial transformations to guide and regularize registration networks, ensuring anatomically plausible deformations. It features a bottleneck autoencoder that computes the principal logarithm of deformation fields via iterative square-root predictions. It creates a linearized latent space that respects diffeomorphic properties and enforces inverse consistency. By integrating a registration network with a diffeomorphic autoencoder, MORPH-LER produces smooth, meaningful deformation fields. The framework offers two main contributions: (1) a data-driven regularization strategy that incorporates population-level anatomical statistics to enhance transformation validity and (2) a linearized latent space that enables compact and interpretable deformation fields for efficient population morphometrics analysis. We validate MORPH-LER across two families of deep learning-based registration networks, demonstrating its ability to produce anatomically accurate, computationally efficient, and statistically meaningful transformations on the OASIS-1 brain imaging dataset. https://github.com/iyerkrithika21/MORPH_LER
@@ 1 Introduction
Deep learning methods address these challenges by learning complex transformations directly from images, offering faster inference and ensuring smooth deformation fields through regularizers like bending energy or $L_{2}-$gradient penalties [Balakrishnan et al.(2019)Balakrishnan, Zhao, Sabuncu, Guttag, and Dalca, Chen et al.(2022)Chen, Frey, He, Segars, Li, and Du]. However, they treat registration as a computational problem, neglecting diffeomorphic properties like inversion and composition, resulting in biologically implausible solutions. Improvements like gradient and transformation inverse consistency enhance diffeomorphic transformations [Tian et al.(2023)Tian, Greer, Vialard, Kwitt, Estépar, Rushmore, Makris, Bouix, and Niethammer]. Yet, existing approaches often predict transformations without deeper insights into underlying anatomical relationships, oversimplifying interpretations.
Nevertheless, the mathematical framework of diffeomorphic transformations (smooth, invertible transformations that preserve the topology of the image, meaning they don’t tear or fold it) defined by the diffeomorphism group $Diff(\mathcal{M})$ (the set of all such transformations), which, mathematically, has the structure of a Fréchet Lie group and manifold provides a rigorous foundation for capturing and analyzing complex anatomical variability. Understanding this framework is essential for ensuring image registration methods produce realistic and meaningful results.
Unlike Euclidean spaces, operations like addition or averaging cannot be directly defined on the diffeomorphism group $Diff(\mathcal{M})$. This means we can’t simply average two diffeomorphic transformations to get a mean transformation in the same way we can average two points in a Euclidean space. Techniques such as Principal Geodesic Analysis (PGA) [Fletcher et al.(2004)Fletcher, Lu, Pizer, and Joshi], Fréchet means [Le and Kume(2000)], and geodesic regression [Fletcher(2011)] attempt to address statistical estimation for Riemannian manifolds (curved spaces where distances and angles can be measured), but require significant adaptation for Fréchet Lie groups. These techniques need to be modified to account for the complex structure of diffeomorphism groups. Therefore, alternative metrics are necessary to model relationships between deformation fields (the spatial displacements that map one image to another) and ensure anatomical relevance. The Log-Euclidean framework [Arsigny et al.(2006)Arsigny, Commowick, Pennec, and Ayache] transforms non-linear diffeomorphism manifolds into linear Lie algebra spaces (vector spaces that approximate the diffeomorphism group locally), enabling efficient computation of transformation distances and statistical analysis. However, iterative processes used in logarithm estimation can be susceptible to initializations, potentially leading to suboptimal results. Standalone deformation analysis models address the limitations of the iterative methods and provide a computationally efficient transformation of diffeomorphic groups into linear Lie algebra. By mapping the problem to a Lie algebra, we can perform calculations more easily. These mathematical foundations provide a rigorous framework for analyzing the complex structure of diffeomorphisms, leading to more reliable and interpretable results.
@@ 2 Related Work
Population-aware registration networks, such as CAE [Bhalodia et al.(2019)Bhalodia, Elhabian, Kavan, and Whitaker], add regularization by limiting deformation fields to low-dimensional manifolds, effectively capturing anatomical variability. However, they often fall short of explicitly enforcing diffeomorphism, essential for interpretability and robustness. CAE’s regularization approach begins with an L2 regularizer, providing initial stability. However, when the autoencoder-based regularizer is introduced, it struggles to preserve anatomical consistency, often resulting in non-diffeomorphic transformations.
Similarly, diffeomorphic autoencoders for LDDMM [Hinkle et al.(2018)Hinkle, Womble, and Yoon] have been proposed for atlas building by encoding velocity field-based transformations into latent spaces for statistical analysis. These methods face computational challenges with large datasets and complex anatomical variations. Additionally, the focus on deformation field reconstruction from latent space representation often sacrifices interpretability, limiting their utility for downstream tasks such as population analysis. The Log Euclidean Diffeomorphic Autoencoder (LEDA) [Iyer et al.(2024)Iyer, Elhabian, and Joshi] framework has combined the advantages of the Log Euclidean framework with a deep learning approach. LEDA estimates the principal logarithms of the deformation field while maintaining inverse consistency in both the deformation and latent spaces. This approach efficiently captures complex anatomical variations while remaining computationally tractable for large-scale analysis. It maps composition in the data space to linearized latent space, enabling vector-based statistics and ensuring robustness while adhering to diffeomorphism group laws.
@@
Figure 1: Proposed Architecture for MORPH-LER
@@ 3 Methods
Background: A diffeomorphism ${\boldsymbol{\phi}}:\mathbb{R}^{D}\to\mathbb{R}^{D}$, is a smooth, invertible mapping with a smooth inverse, defined as ${\boldsymbol{\phi}}(\mathbf{x})=\mathbf{x}+\mathbf{u}(\mathbf{x})$, where $\mathbf{u}(\mathbf{x})$ is the displacement field. Diffeomorphic transformations can be constructed by composing multiple small deformations: ${\boldsymbol{\phi}}(\mathbf{x})=(\mathbf{x}+\epsilon\mathbf{v}_{1}(\mathbf{x}))\circ(\mathbf{x}+\epsilon\mathbf{v}_{2}(\mathbf{x}))\circ\cdots\circ(\mathbf{x}+\epsilon\mathbf{v}_{n}(\mathbf{x}))$. The space of diffeomorphisms $\mathrm{Diff}(\mathcal{M})$ forms a Lie group under composition, with the identity map serving as the group’s identity element; i.e., ${\boldsymbol{\phi}}\circ\mathbf{id}=\mathbf{id}\circ{\boldsymbol{\phi}}={\boldsymbol{\phi}},\forall{\boldsymbol{\phi}}\in\mathrm{Diff}(\mathcal{M})$. The associated Lie algebra consists of smooth vector fields $\mathbf{v}(\mathbf{x}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, which generate diffeomorphisms through their flows, described by one-parameter subgroups $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$. The exponential map $\exp(\mathbf{v})$ connects the Lie algebra to the Lie group, while the logarithm map $\log({\boldsymbol{\phi}})$ serves as its local inverse. To compute the logarithm map, a non-linear inverse scaling and square rooting algorithm [Arsigny et al.(2006)Arsigny, Commowick, Pennec, and Ayache] is used:
@@
Table 1: Registration Performance Metrics: Top-performing model and next-best model.
@@
Figure 2: (A) MORPH-LER Square Root Estimation: GradICON (left) and TranMorph (right) as the primary registration network. (B) Validation of Small Deformation Field Assumption. (C) Validation of Latent Inverse Consistency.
@@ 4 Results
Figure 3: MORPH-LER PCA Modes of Logarithm Maps and Latent Representations. Red arrows highlight key structural shape changes for each PCA mode. We display the deformation field and apply it to a randomly chosen image to illustrate modes of variation in the population of the deformation field.
Figure 4: Qualitative Registration Results: Four-label brain segmentation (cortex, gray matter, sub-cortical gray matter, CSF) used to assess registration quality. The table shows Dice scores between warped and target segmentations for the visualized sample, with color-coding matching segmentation regions. Top (bold) and runner-up (underlined) models are highlighted for the visualized sample.
@@
We use 2D coronal slices from the OASIS-1 dataset [Marcus et al.(2007)Marcus, Wang, Parker, Csernansky, Morris, and Buckner], which includes brain MRIs from 100 subjects (60 with Alzheimer’s). We evaluate our proposed regularizer against several baselines: CAE [Bhalodia et al.(2019)Bhalodia, Elhabian, Kavan, and Whitaker], an autoencoder-based method, and the original GradICON [Tian et al.(2023)Tian, Greer, Vialard, Kwitt, Estépar, Rushmore, Makris, Bouix, and Niethammer] and TransMorph [Chen et al.(2022)Chen, Frey, He, Segars, Li, and Du] without LEDA. GradICON penalizes deviations from the Jacobian of the inverse consistency constraint, while TransMorph combines a Swin-Transformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] encoder and a convolutional decoder for volumetric medical image registration. We utilize GradICON and TransMorph as primary registration networks with LEDA as the secondary network, naming them M-GradICON and M-TransMorph. Table 1 compares TransMorph vs. M-TransMorph, GradICON vs. M-GradICON, and CAE as a standalone regularizer. Within each subgroup, we analyze trade-offs between registration accuracy and topological preservation. M-TransMorph improves diffeomorphic properties by reducing negative Jacobian pixels, but slightly impacts segmentation performance. M-GradICON achieves higher Dice scores than GradICON while maintaining low negative Jacobian pixels, balancing accuracy and topology preservation. Compared to CAE, LEDA-based regularization maintains similar Dice scores while significantly reducing negative Jacobian pixels, demonstrating its superiority in preserving anatomical topology without compromising accuracy. The improvements in metrics achieved by MORPH-LER variants were statistically significant, determined by paired t-tests $(p<0.05)$ across all model comparisons.
@@
Figure 4 showcases a simplified 4-label brain segmentation used to assess registration quality. The segmentation comprises cortex, gray matter, subcortical gray matter, and cerebrospinal fluid, providing a concise yet informative representation of key brain structures. We apply learned deformation fields to these segmentation labels to evaluate registration performance and compute Dice scores between the warped and target segmentations. These results corroborate the quantitative findings. Notably, our proposed M-GradICON model significantly improves over its baseline counterpart. This underscores the effectiveness of the secondary network. The performance of M-GradICON is especially noticeable in the smaller anatomical regions, such as subcortical gray matter and cerebrospinal fluid. The quantitative and qualitative results show that the CAE regularization strategy with a simple autoencoder is ineffective for complex shapes, lacks diffeomorphic properties, and produces anatomically inconsistent transformations. Appendix Figure 6 shows more examples of CAE on toy dataset.
@@
Figure 2.A illustrates the square root estimations $\phi^{-m}$ of MORPH-LER variants and demonstrates a progressive warping of the source image to align with the target. M-GradICON exhibits superior performance over M-TransMorph, featuring smoother deformation grids, better alignment of warped images, and smoother Jacobian maps, indicating diffeomorphic superiority. To validate the small deformation field assumption [^1] Small deformation field assumption $u_{AB}(\mathbf{x})=u_{BA}(\mathbf{x})$ reflects symmetry of infinitesimal displacements, ensuring consistent and reversible transformations., we tested the logarithm map consistency by utilizing the $2^{6}$-th root estimation, we systematically negated and composed forward and inverse displacement fields. Figure 2.C demonstrates that both methods satisfy this fundamental assumption, validating the accuracy of logarithm maps.
@@
Furthermore, we extended this validation to the model’s latent space by negating the latent representations of forward and inverse displacement fields. As illustrated in Figure 2.B, the MORPH-LER accurately decodes these negated latent representations into their corresponding inverse fields, providing compelling evidence of inverse consistency in the latent space. PCA analysis of the MORPH-LER’s latent space, depicted in Figure 3, reveals modes of variation that align closely with logarithm map PCA results. Red arrows highlight key structural shape changes for each PCA mode. These modes capture clinically consistent changes, such as ventricular expansion and hippocampal atrophy, while maintaining smooth, structured transitions. This latent space representation not only supports accurate reconstruction of deformations but also enables intuitive exploration of clinically meaningful variations, highlighting the MORPH-LER’s potential for generating realistic deformations, interpolating between anatomical states, and providing valuable insights for understanding disease progression.
@@
To create a population-level representation that describes the average shape, and structure, of a population of anatomical objects [Joshi et al.(2004)Joshi, Davis, Jomier, and Gerig] we propose an efficient atlas estimation approach that leverages the trained MORPH-LER’s linearized latent space, which adheres to Lie group action laws. The algorithm begins by randomly selecting an initial image $\mathbf{A}^{(0)}$ as the starting atlas. For each image $\mathbf{I}_{i}$ in the dataset, the algorithm computes bidirectional transformations (${\boldsymbol{\phi}}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$ and ${\boldsymbol{\phi}}_{\mathbf{I}_{i}\mathbf{A}^{(k)}}$) between the current atlas $\mathbf{A}^{(k)}$ and the image $\mathbf{I}_{i}$. The latent representations of these transformations, $\mathbf{z}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$ and $\mathbf{z}_{\mathbf{I}_{i}\mathbf{A}^{(k)}}$, are extracted using the LEDA module. At each iteration, the latent representations corresponding to the transformations from the atlas to all images, $\mathbf{z}_{\mathbf{A}^{(k)}\mathbf{I}_{i}}$, are averaged across all $i$ to compute a mean representation, $\mathbf{z}^{k}$. This negated mean representation is decoded to obtain the atlas to image mean deformation field $\overline{{\boldsymbol{\phi}}}^{k}$, which is then used to update the atlas $\mathbf{A}^{(k)}$ pulling it towards the true mean. The process is repeated until the atlas converges to a stable solution, as determined by minimal changes across iterations. Additional details are provided in the Appendix. To evaluate the robustness of this method, we perform multiple estimations, each initialized with a different atlas. The initial atlases are chosen randomly with respect to two distinct age groups: above 45 years and below 45 years. As shown in Figure 5, the final estimated atlas remains an unbiased estimate regardless of the initialization, demonstrating the robustness of the approach. Unlike a naive pixel-wise average that may introduce artifacts, the proposed approach ensures consistent geometric alignment across the dataset. This alignment is crucial for studying population variability and enables effective downstream statistical analysis, ensuring that the atlas remains biologically meaningful and robust for comparative studies.
Figure 5: Estimated atlas using the proposed algorithm compared with pixel-wise average atlas.